Q: If a, b and y are positive real numbers with such that none of them is equal to 1 and b^2a+6 = y^2, which of these must be true?
A) y = b^a+3
B) 2a + 6 = 2
C) b = y
D) b = 2
E) b(2a + 6) = 2y
Explain too plz :)

Answer by Boreal(15235)

b^2a+6 = y^2,
y= sqrt (b^(2a +6)), is y=b^(a+3) and that works.
let a=3 and b=2.
2^(12)=y^2
y=2^6=64
y=b^(a+3)=2^6=64
y=sqrt(b^2a+6)=[b^(2a+6)]^(1/2), and that is b^(a+3) ANSWER
---------------------
2a+6=2.
If b=y, this is true. If it is b^(2a+6)=b^2, then it works in all cases.
Suppose, however, b=2 and a=3 Then 2^(12)=y^2, and y^2=4096 and y=64.
---------------------
b(2a+6)=2y is a misunderstanding of raising to a power. b^2a+6 is raising to a power, not multiplying.
--------------------
b=2. It can be, but they are all positive real numbers except 1.
-------------------
b=y only if 2a+6=2. Since they can be positive real numbers except 1, that isn't "must"
--------------------